# Overlap reduction functions for pulsar timing arrays and astrometry

### v1

@@ III.1 Auto-correlation of redshift responses
$z(t|\hat{\Omega})=\frac{\hat{n}_{a}^{i}\hat{n}_{a}^{j}\Delta h_{ij}(t,\hat{\Omega})}{2(1+\hat{\Omega}\cdot\hat{n}_{a})},$ (17)
@@
$z_{f}(\hat{n})$ $=\sum_{\ell m}Y_{\ell m}(\theta,\phi)z_{f,lm}.$ (18)
Note that $z_{f}$ is a Fourier mode and can be complex.
@@
$=\sum_{p=t,v}\sum_{X}i^{X}\left[I^{p}_{0}\sum_{L,M}c^{p,X}_{LM}\Gamma^{p,X,zz}_{LM}(\hat{n}_{a},\hat{n}_{b})\right],$ (19)
@@ III.1.1 Spin-2 GWs
$z_{f}(\hat{n}|\hat{k})$ $=\frac{1}{2}\left[\tilde{h}^{+}(f,\hat{k})(1-\cos\theta)\cos(2\phi)+\tilde{h}^{\times}(f,\hat{k})(1-\cos\theta)\sin(2\phi)\right]$
@@
$z_{f}(\hat{n}|\hat{\Omega})=\sum_{\ell}\sum_{mm^{\prime}}Y_{\ell m}(\hat{\Omega})D^{(\ell)}_{mm^{\prime}}(\hat{\Omega})\tilde{z}_{f,\ell m^{\prime}}(\hat{k}),$ (25)
where $D^{(\ell)}_{mm^{\prime}}(\hat{\Omega})=D^{(\ell)}_{mm^{\prime}}(\phi,\theta,0)$ with $\hat{\Omega}=(\theta,\phi)$. See also Appendix A for the properties of Wigner $D$-matrix. From this, we obtain
$\tilde{z}_{f,\ell m}(\hat{\Omega})$ $=\int\mathrm{d}\hat{n}\,Y^{*}_{\ell m}(\hat{n})z_{f}(\hat{n}|\hat{\Omega})$
@@
$=\frac{z^{t}_{\ell}z^{t}_{\ell^{\prime}}}{2}I^{t}_{0}\sum_{LM}\left[c^{t,I}_{LM}\left(\,{}^{(0,2,2)}W^{LM}_{\ell\ell^{\prime}mm^{\prime}}+\,^{(0,-2,-2)}W^{LM}_{\ell\ell^{\prime}mm^{\prime}}\right)+c^{t,V}_{LM}\left(\,{}^{(0,2,2)}W^{LM}_{\ell\ell^{\prime}mm^{\prime}}-\,^{(0,-2,-2)}W^{LM}_{\ell\ell^{\prime}mm^{\prime}}\right)\right.$
$\quad\qquad\qquad\qquad\left.+c^{t,+}_{LM}\,{}^{(4,2,-2)}W^{LM}_{\ell\ell^{\prime}mm^{\prime}}+c^{t,-}_{LM}\,{}^{(-4,-2,2)}W^{LM}_{\ell\ell^{\prime}mm^{\prime}}\right].$ (27)
@@
${z^{*}_{f,\ell m}z_{f,\ell^{\prime}m^{\prime}}}\right>$ $=2z^{t}_{\ell}z^{t}_{\ell^{\prime}}I^{t}_{0}\sum_{LM}(-1)^{m+L}\sqrt{\pi}\Braket{\ell(-m)\ell^{\prime}m^{\prime}}{LM}$
×$\qquad\times\left\{\left[c^{t,I}_{LM}X^{L}_{\ell\ell^{\prime}}-c^{t,V}_{LM}(1-X^{L}_{\ell\ell^{\prime}})\right]\begin{pmatrix}\ell&\ell^{\prime}&L\\ -2&2&0\\ \end{pmatrix}+\left[c^{t,E}_{LM}X^{L}_{\ell\ell^{\prime}}+ic^{t,B}_{LM}(1-X^{L}_{\ell\ell^{\prime}})\right]\begin{pmatrix}\ell&\ell^{\prime}&L\\ 2&2&-4\\ \end{pmatrix}\right\},$ (33)
@@
In these coordinates, the ORF only depends on $\theta$. Figure 1 shows the $\theta$ dependence of $\Gamma^{t,I,zz}_{LM}$ with different $\ell_{\text{max}}$, where we also show the exact analytical results in Ref. [35]. From this figure, we can see that $\ell_{\text{max}}=10$ is a good value for the convergence for spin-2 GW case.
@@ III.1.2 Spin-1 GWs
$z(\hat{n}|\hat{k})$ $=\frac{1}{2}\left[\tilde{h}^{x}\frac{\sin 2\theta\cos\phi}{1+\cos\theta}+\tilde{h}^{y}\frac{\sin 2\theta\sin\phi}{1+\cos\theta}\right]$
@@
Similar to the spin-2 GW case, we can find $z_{f,\ell m}=0$ in $\ell<1$, which leads to $\ell_{\text{min}}=1$ in Eq. (20) for the spin-1 GWs. We here generalize this to the case of GWs propagating in $\hat{\Omega}$:
$\tilde{z}_{f,\ell m}(\hat{\Omega})$ $=\int\mathrm{d}\hat{n}\,Y^{*}_{\ell m}(\hat{n})z_{f}(\hat{n}|\hat{\Omega})$
@@
$=\frac{z^{v}_{\ell}z^{v}_{\ell^{\prime}}}{2}I^{v}_{0}\sum_{LM}\left[c^{v,I}_{LM}\left(\,{}^{(0,1,1)}W^{LM}_{\ell\ell^{\prime}mm^{\prime}}+\,^{(0,-1,-1)}W^{LM}_{\ell\ell^{\prime}mm^{\prime}}\right)+c^{v,V}_{LM}\left(\,{}^{(0,1,1)}W^{LM}_{\ell\ell^{\prime}mm^{\prime}}-\,^{(0,-1,-1)}W^{LM}_{\ell\ell^{\prime}mm^{\prime}}\right)\right.$
$\quad\qquad\qquad\qquad\left.-c^{v,+}_{LM}\,{}^{(2,1,-1)}W^{LM}_{\ell\ell^{\prime}mm^{\prime}}-c^{v,-}_{LM}\,{}^{(-2,-1,1)}W^{LM}_{\ell\ell^{\prime}mm^{\prime}}\right]$
$=2z^{v}_{\ell}z^{v}_{\ell^{\prime}}I^{v}_{0}\sum_{LM}(-1)^{m+L}\sqrt{\pi}\Braket{\ell(-m)\ell^{\prime}m^{\prime}}{LM}$
×$\qquad\times\left\{\left[-c^{v,I}_{LM}X^{L}_{\ell\ell^{\prime}}+c^{v,V}_{LM}(1-X^{L}_{\ell\ell^{\prime}})\right]\begin{pmatrix}\ell&\ell^{\prime}&L\\ -1&1&0\\ \end{pmatrix}+\left[c^{v,E}_{LM}X^{L}_{\ell\ell^{\prime}}+ic^{v,B}_{LM}(1-X^{L}_{\ell\ell^{\prime}})\right]\begin{pmatrix}\ell&\ell^{\prime}&L\\ 1&1&-2\\ \end{pmatrix}\right\},$ (39)
Substituting this into Eq. (19), we obtain the coefficient $F^{L}_{\ell\ell^{\prime}}$, summarized in the right column of Table 1. Figure 2 shows $\Gamma^{v,I,zz}_{LM}(\theta)$ with different $\ell_{\text{max}}$ in the coordinates of Eq. (34). In the figure, We also show the exact analytical results in Ref. [35]. The convergence with respect to $\ell_{\text{max}}$ is slower than the spin-2 GW case (Fig. 1). This is because $z^{v}_{\ell}$ decays slower than $z^{t}_{\ell}$ in $\ell\gg 1$ ($z^{v}_{\ell}\propto\ell^{-1/2}$ and $z^{t}_{\ell}\propto\ell^{-3/2}$ in $\ell\gg 1$).
@@ III.2 Auto-correlation of deflection fields
$\delta n^{i}(t|\hat{n})=\frac{(\hat{n}^{i}+\hat{\Omega}^{i})\hat{n}^{j}\hat{n}^{k}h_{jk}(t,\hat{\Omega},\vec{0})}{2(1+\hat{n}\cdot\hat{\Omega})}-\frac{1}{2}\hat{n}^{j}h_{ij}(t,\hat{\Omega},\vec{0}),$ (40)
where the origin $\vec{0}$ corresponds to the observer position. In the following, we raise or lower the spatial indices with the Kronecker delta, e.g. $\delta n^{i}=\delta^{ij}\delta n_{j}$. We expand the deflection vector $\delta\vec{n}$ as
$\delta n_{i}(\hat{n})$ $=\sum_{\ell m}\left[E_{\ell m}Y^{E}_{(\ell m)i}+B_{\ell m}Y^{B}_{(\ell m)i}\right],$ (41)
@@
From Eq. (40), we can also define a scalar $\bar{E}(t|\hat{n})$ and pseudoscalar $\bar{B}(t|\hat{n})$ induced by GWs propagating in $\hat{\Omega}$ as
$\bar{E}(t|\hat{n})$ $=-\frac{1}{2}\text{Tr}\,h(t,\hat{\Omega},\vec{0})+\frac{(\hat{n}^{i}+\hat{\Omega}^{i})\hat{n}^{j}h_{ij}(t,\hat{\Omega},\vec{0})}{1+\hat{\Omega}\cdot\hat{n}},$ (49)
$\bar{B}(t|\hat{n})$ $=\epsilon^{ijk}\frac{\hat{\Omega}_{i}\hat{n}^{l}\hat{n}_{k}h_{jl}(t,\hat{\Omega},\vec{0})}{1+\hat{\Omega}\cdot\hat{n}},$ (50)
@@
$=\sum_{p=t,v}\sum_{X}\sum_{S,T}i^{X}\left[I^{p}_{0}\sum_{L,M}c^{p,X}_{LM}\Gamma^{p,X,ST}_{LM,ij}(\hat{n}_{a},\hat{n}_{b})\right],$ (51)
@@ III.2.1 Spin-2 GWs
$\bar{E}_{f}(\hat{n}|\hat{k})$ $=\tilde{h}_{+}(1-\cos\theta)\cos(2\phi)+\tilde{h}_{\times}(1-\cos\theta)\sin(2\phi)$
@@
$\bar{B}_{f}(\hat{n}|\hat{k})$ $=\tilde{h}_{+}(1-\cos\theta)\sin 2\phi\,-\tilde{h}_{\times}(1-\cos\theta)\cos 2\phi$
@@
$\tilde{E}_{f,\ell m}(\hat{k})$ $=\frac{1}{\sqrt{\ell(\ell+1)}}\int\mathrm{d}^{2}\hat{n}\,Y^{*}_{\ell m}(\hat{n})\bar{E}_{f}(\hat{n}|\hat{k})$
@@
$\tilde{B}_{f,\ell m}(\hat{k})$ $=\frac{1}{\sqrt{\ell(\ell+1)}}\int\mathrm{d}^{2}\hat{n}\,Y^{*}_{\ell m}(\hat{n})\bar{B}_{f}(\hat{n}|\hat{k})$
@@
$=2E^{t}_{\ell}B^{t}_{\ell^{\prime}}I^{t}_{0}\sum_{LM}(-1)^{m+L}\sqrt{\pi}\Braket{\ell(-m)\ell^{\prime}m^{\prime}}{LM}$
×$\qquad\times\left\{\left[-c^{t,I}_{LM}(1-X^{L}_{\ell\ell^{\prime}})+c^{t,V}_{LM}X^{L}_{\ell\ell^{\prime}}\right]\begin{pmatrix}\ell&\ell^{\prime}&L\\ -2&2&0\\ \end{pmatrix}-\left[c^{t,E}_{LM}(1-X^{L}_{\ell\ell^{\prime}})+ic^{t,B}_{LM}X^{L}_{\ell\ell^{\prime}}\right]\begin{pmatrix}\ell&\ell^{\prime}&L\\ 2&2&-4\\ \end{pmatrix}\right\},$ (62)
@@
$=2B^{t*}_{\ell}E^{t}_{\ell^{\prime}}I^{t}_{0}\sum_{LM}(-1)^{m+L}\sqrt{\pi}\Braket{\ell(-m)\ell^{\prime}m^{\prime}}{LM}$
×$\qquad\times\left\{\left[-c^{t,I}_{LM}(1-X^{L}_{\ell\ell^{\prime}})+c^{t,V}_{LM}X^{L}_{\ell\ell^{\prime}}\right]\begin{pmatrix}\ell&\ell^{\prime}&L\\ -2&2&0\\ \end{pmatrix}+\left[c^{t,E}_{LM}(1-X^{L}_{\ell\ell^{\prime}})+ic^{t,B}_{LM}X^{L}_{\ell\ell^{\prime}}\right]\begin{pmatrix}\ell&\ell^{\prime}&L\\ 2&2&-4\\ \end{pmatrix}\right\},$ (63)
@@
$=2B^{t*}_{\ell}B^{t}_{\ell^{\prime}}I^{t}_{0}\sum_{LM}(-1)^{m+L}\sqrt{\pi}\Braket{\ell(-m)\ell^{\prime}m^{\prime}}{LM}$
×$\qquad\times\left\{\left[c^{t,I}_{LM}X^{L}_{\ell\ell^{\prime}}-c^{t,V}_{LM}(1-X^{L}_{\ell\ell^{\prime}})\right]\begin{pmatrix}\ell&\ell^{\prime}&L\\ -2&2&0\\ \end{pmatrix}-\left[c^{t,E}_{LM}X^{L}_{\ell\ell^{\prime}}+ic^{t,B}_{LM}(1-X^{L}_{\ell\ell^{\prime}})\right]\begin{pmatrix}\ell&\ell^{\prime}&L\\ 2&2&-4\\ \end{pmatrix}\right\},$ (64)
@@
and the left column in Tables 2 for $F^{L,t,X,EB}_{\ell\ell^{\prime}}$, 3 for $F^{L,t,X,BE}_{\ell\ell^{\prime}}$, and 4 for $F^{L,t,X,BB}_{\ell\ell^{\prime}}$. To get the concrete plots of the ORFs, we similarly take the coordinates given in Eq. (34). Strictly speaking, in that choice of coordinates, $Y^{E}_{(\ell m)\phi}(\hat{n}_{a})$ and $Y^{B}_{(\ell m)\theta}(\hat{n}_{a})$ are not well defined because of $\sin\theta$ in the denominator, which is from the singularity of the coordinates. To avoid this issue, we define $Y^{E}_{(\ell m)\phi}(\hat{n}_{a})$ and $Y^{B}_{(\ell m)\theta}(\hat{n}_{a})$ with $\hat{n}_{a}=\hat{n}_{b}|_{\theta\rightarrow+0}$ in the coordinates. Practically, we take $\hat{n}_{a}=(\sin\theta_{\epsilon},0,\cos\theta_{\epsilon})$ with $\theta_{\epsilon}=10^{-5}$ in the numerical calculation for Figs. 3-6. Then, Figure 3 shows the $\theta$-dependence of $\Gamma^{t,I,EB}_{LM}$. In the figure, we can see $\Gamma^{t,I,EB}_{00}=0$. This physically means that the parity-breaking signal ($EB$-correlation) cannot be produced in the parity-conserving background (isotropic and unpolarized background).

### v2

@@ III.1 Auto-correlation of redshift responses
$z(t,\hat{n}_{a},\hat{\Omega})=\frac{\hat{n}_{a}^{i}\hat{n}_{a}^{j}\Delta h_{ij}(t,\hat{\Omega})}{2(1+\hat{\Omega}\cdot\hat{n}_{a})},$ (17)
@@
$z_{f}(\hat{n})$ $=\sum_{\ell m}Y_{\ell m}(\theta,\phi)z_{f,\ell m}.$ (18)
Note that $z_{f}$ is a Fourier mode and can be complex and the subscript $f$ is the shorthand notation for the Fourier mode argument, e.g. $z_{f}=z(f)$, throughout this work.
@@
$=\sum_{p=t,v}\sum_{X}i^{X}\left[I^{p}_{0}(f)\sum_{L,M}c^{p,X}_{LM}(f)\Gamma^{p,X,zz}_{LM}(\hat{n}_{a},\hat{n}_{b})\right],$ (19)
@@ III.1.1 Spin-2 GWs
$z_{f}(\hat{n},\hat{k})$ $=\frac{1}{2}\left[\tilde{h}^{+}(f,\hat{k})(1-\cos\theta)\cos(2\phi)+\tilde{h}^{\times}(f,\hat{k})(1-\cos\theta)\sin(2\phi)\right]$
@@
$z_{f}(\hat{n},\hat{\Omega})=\sum_{\ell}\sum_{mm^{\prime}}Y_{\ell m}(\hat{n})D^{(\ell)}_{mm^{\prime}}(\hat{\Omega})\tilde{z}_{f,\ell m^{\prime}}(\hat{\Omega}),$ (25)
where $D^{(\ell)}_{mm^{\prime}}(\hat{\Omega})=D^{(\ell)}_{mm^{\prime}}(\phi_{\Omega},\theta_{\Omega},0)$ with $\hat{\Omega}=(\theta_{\Omega},\phi_{\Omega})$. See also Appendix A for the properties of Wigner $D$-matrix. From this, we obtain
$\tilde{z}_{f,\ell m}(\hat{\Omega})$ $=\int\mathrm{d}\hat{n}\,Y^{*}_{\ell m}(\hat{n})z_{f}(\hat{n},\hat{\Omega})$
@@
$=\frac{z^{t}_{\ell}z^{t}_{\ell^{\prime}}}{2}I^{t}_{0}(f)\sum_{LM}\left[c^{t,I}_{LM}(f)\left(\,{}^{(0,2,2)}W^{LM}_{\ell\ell^{\prime}mm^{\prime}}+\,^{(0,-2,-2)}W^{LM}_{\ell\ell^{\prime}mm^{\prime}}\right)+c^{t,V}_{LM}(f)\left(\,{}^{(0,2,2)}W^{LM}_{\ell\ell^{\prime}mm^{\prime}}-\,^{(0,-2,-2)}W^{LM}_{\ell\ell^{\prime}mm^{\prime}}\right)\right.$
$\quad\qquad\qquad\qquad\left.+c^{t,+}_{LM}(f)\,^{(4,2,-2)}W^{LM}_{\ell\ell^{\prime}mm^{\prime}}+c^{t,-}_{LM}(f)\,^{(-4,-2,2)}W^{LM}_{\ell\ell^{\prime}mm^{\prime}}\right].$ (27)
@@
${z^{*}_{f,\ell m}z_{f,\ell^{\prime}m^{\prime}}}\right>$ $=2z^{t}_{\ell}z^{t}_{\ell^{\prime}}I^{t}_{0}(f)\sum_{LM}(-1)^{m+L}\sqrt{\pi}\Braket{\ell(-m)\ell^{\prime}m^{\prime}}{LM}$
×$\quad\times\left\{\left[c^{t,I}_{LM}(f)X^{L}_{\ell\ell^{\prime}}-c^{t,V}_{LM}(f)(1-X^{L}_{\ell\ell^{\prime}})\right]\begin{pmatrix}\ell&\ell^{\prime}&L\\ -2&2&0\\ \end{pmatrix}+\left[c^{t,E}_{LM}(f)X^{L}_{\ell\ell^{\prime}}+ic^{t,B}_{LM}(f)(1-X^{L}_{\ell\ell^{\prime}})\right]\begin{pmatrix}\ell&\ell^{\prime}&L\\ 2&2&-4\\ \end{pmatrix}\right\},$ (33)
@@
In these coordinates, the ORF only depends on $\theta$. Figure 1 shows the $\theta$ dependence of $\Gamma^{t,I,zz}_{LM}$ with different $\ell_{\text{max}}$, where we also show the exact analytical results in Ref. [35]. Figure 2 shows the $\ell_{\text{max}}$ dependence of the deviation from the exact value at $\theta=0$, where the deviation becomes the largest. From these figures, we can see that $\ell_{\text{max}}=10$ is a good value for the convergence for spin-2 GW case. For example, the deviation from the exact value at $\theta=0$ in $L=1$ and $M=0$ is $\sim 5\%$ when $\ell_{\text{max}}=10$. In general, the value of $\ell_{\text{max}}$ must be tuned to balance the computational time and the required precision for the comparison with real data, while $\mathcal{O}(1)\,\%$ precision would be sufficient for near-future experiments.
@@
Figure 2: The $\ell_{\text{max}}$ dependence of $\Gamma^{t,I,zz}_{LM}(\theta=0)$ normalized by the exact analytical results obtained in Ref. [35].
@@ III.1.2 Spin-1 GWs
$z(\hat{n},\hat{k})$ $=\frac{1}{2}\left[\tilde{h}^{x}\frac{\sin 2\theta\cos\phi}{1+\cos\theta}+\tilde{h}^{y}\frac{\sin 2\theta\sin\phi}{1+\cos\theta}\right]$
@@
Similar to the spin-2 GW case, we can find $z_{f,\ell m}=0$ in $\ell<1$, which leads to $\ell_{\text{min}}=1$ in Eq. (20) for the spin-1 GWs. We here note that the relative sign between the two terms in Eq. (36) is different from the spin-2 GW case (Eq. (23)). Technically, this difference comes from the overall sign difference for $Y_{\ell 1}(\theta,0)$ and $Y_{\ell,-1}(\theta,0)$. We here generalize this to the case of GWs propagating in $\hat{\Omega}$:
$\tilde{z}_{f,\ell m}(\hat{\Omega})$ $=\int\mathrm{d}\hat{n}\,Y^{*}_{\ell m}(\hat{n})z_{f}(\hat{n},\hat{\Omega})$
@@
$=\frac{z^{v}_{\ell}z^{v}_{\ell^{\prime}}}{2}I^{v}_{0}(f)\sum_{LM}\left[c^{v,I}_{LM}(f)\left(\,{}^{(0,1,1)}W^{LM}_{\ell\ell^{\prime}mm^{\prime}}+\,^{(0,-1,-1)}W^{LM}_{\ell\ell^{\prime}mm^{\prime}}\right)+c^{v,V}_{LM}(f)\left(\,{}^{(0,1,1)}W^{LM}_{\ell\ell^{\prime}mm^{\prime}}-\,^{(0,-1,-1)}W^{LM}_{\ell\ell^{\prime}mm^{\prime}}\right)\right.$
$\quad\qquad\qquad\qquad\left.-c^{v,+}_{LM}(f)\,^{(2,1,-1)}W^{LM}_{\ell\ell^{\prime}mm^{\prime}}-c^{v,-}_{LM}(f)\,^{(-2,-1,1)}W^{LM}_{\ell\ell^{\prime}mm^{\prime}}\right]$
$=2z^{v}_{\ell}z^{v}_{\ell^{\prime}}I^{v}_{0}(f)\sum_{LM}(-1)^{m+L}\sqrt{\pi}\Braket{\ell(-m)\ell^{\prime}m^{\prime}}{LM}$
×$\qquad\times\left\{\left[-c^{v,I}_{LM}(f)X^{L}_{\ell\ell^{\prime}}+c^{v,V}_{LM}(f)(1-X^{L}_{\ell\ell^{\prime}})\right]\begin{pmatrix}\ell&\ell^{\prime}&L\\ -1&1&0\\ \end{pmatrix}+\left[c^{v,E}_{LM}(f)X^{L}_{\ell\ell^{\prime}}+ic^{v,B}_{LM}(f)(1-X^{L}_{\ell\ell^{\prime}})\right]\begin{pmatrix}\ell&\ell^{\prime}&L\\ 1&1&-2\\ \end{pmatrix}\right\}.$ (39)
Substituting this into Eq. (19), we obtain the coefficient $F^{L}_{\ell\ell^{\prime}}$, summarized in the right column of Table 1. We here note that the sign difference from Eq. (33) comes from the combination of the relative sign in Eq. (36) and the overall additional minus sign in Eq. (39), which originates from the odd integer $s$ in $D^{(\ell)}_{m,s}$ in Eq. (38). The same sign difference between the spin-2 and the spin-1 GW cases appears in the following subsections (see Tables 2-4) and can be explained in the same way. Figure 3 shows $\Gamma^{v,I,zz}_{LM}(\theta)$ with different $\ell_{\text{max}}$ in the coordinates of Eq. (34). In the figure, We also show the exact analytical results in Ref. [35]. The convergence with respect to $\ell_{\text{max}}$ is slower than the spin-2 GW case (Fig. 1). This is because $z^{v}_{\ell}$ decays slower than $z^{t}_{\ell}$ in $\ell\gg 1$ ($z^{v}_{\ell}\propto\ell^{-1/2}$ and $z^{t}_{\ell}\propto\ell^{-3/2}$ in $\ell\gg 1$).
@@ III.2 Auto-correlation of deflection fields
$\delta n^{i}(t,\hat{n},\hat{\Omega})=\frac{(\hat{n}^{i}+\hat{\Omega}^{i})\hat{n}^{j}\hat{n}^{k}h_{jk}(t,\hat{\Omega},\vec{0})}{2(1+\hat{n}\cdot\hat{\Omega})}-\frac{1}{2}\hat{n}^{j}h_{ij}(t,\hat{\Omega},\vec{0}),$ (40)
where the origin $\vec{0}$ corresponds to the observer position. In the following, we raise or lower the spatial indices with the Kronecker delta, e.g. $\delta n^{i}=\delta^{ij}\delta n_{j}$. Figure 4 visualizes the angles and the deflection vector field. We expand the deflection vector $\delta\vec{n}$ as
$\delta n_{i}(t,\hat{n},\hat{\Omega})$ $=\sum_{\ell m}\left[E_{\ell m}(t,\hat{\Omega})Y^{E}_{(\ell m)i}(\hat{n})+B_{\ell m}(t,\hat{\Omega})Y^{B}_{(\ell m)i}(\hat{n})\right],$ (41)
@@
From Eq. (40), we can also define a scalar $\bar{E}(t,\hat{n})$ and pseudoscalar $\bar{B}(t,\hat{n})$ induced by GWs propagating in $\hat{\Omega}$ as
$\bar{E}(t,\hat{n},\hat{\Omega})$ $=-\frac{1}{2}\text{Tr}\,h(t,\hat{\Omega},\vec{0})+\frac{(\hat{n}^{i}+\hat{\Omega}^{i})\hat{n}^{j}h_{ij}(t,\hat{\Omega},\vec{0})}{1+\hat{\Omega}\cdot\hat{n}},$ (49)
$\bar{B}(t,\hat{n},\hat{\Omega})$ $=\epsilon^{ijk}\frac{\hat{\Omega}_{i}\hat{n}^{l}\hat{n}_{k}h_{jl}(t,\hat{\Omega},\vec{0})}{1+\hat{\Omega}\cdot\hat{n}},$ (50)
@@
$=\sum_{p=t,v}\sum_{X}\sum_{S,T}i^{X}\left[I^{p}_{0}(f)\sum_{L,M}c^{p,X}_{LM}(f)\Gamma^{p,X,ST}_{LM,ij}(\hat{n}_{a},\hat{n}_{b})\right],$ (51)
@@ III.2.1 Spin-2 GWs
$\bar{E}_{f}(\hat{n},\hat{k})$ $=\tilde{h}_{+}(1-\cos\theta)\cos(2\phi)+\tilde{h}_{\times}(1-\cos\theta)\sin(2\phi)$
@@
$\bar{B}_{f}(\hat{n},\hat{k})$ $=\tilde{h}_{+}(1-\cos\theta)\sin 2\phi\,-\tilde{h}_{\times}(1-\cos\theta)\cos 2\phi$
@@
$\tilde{E}_{f,\ell m}(\hat{k})$ $=\frac{1}{\sqrt{\ell(\ell+1)}}\int\mathrm{d}^{2}\hat{n}\,Y^{*}_{\ell m}(\hat{n})\bar{E}_{f}(\hat{n},\hat{k})$
@@
$\tilde{B}_{f,\ell m}(\hat{k})$ $=\frac{1}{\sqrt{\ell(\ell+1)}}\int\mathrm{d}^{2}\hat{n}\,Y^{*}_{\ell m}(\hat{n})\bar{B}_{f}(\hat{n},\hat{k})$
@@
$=2E^{t}_{\ell}B^{t}_{\ell^{\prime}}I^{t}_{0}(f)\sum_{LM}(-1)^{m+L}\sqrt{\pi}\Braket{\ell(-m)\ell^{\prime}m^{\prime}}{LM}$
×$\qquad\times\left\{\left[-c^{t,I}_{LM}(f)(1-X^{L}_{\ell\ell^{\prime}})+c^{t,V}_{LM}(f)X^{L}_{\ell\ell^{\prime}}\right]\begin{pmatrix}\ell&\ell^{\prime}&L\\ -2&2&0\\ \end{pmatrix}-\left[c^{t,E}_{LM}(f)(1-X^{L}_{\ell\ell^{\prime}})+ic^{t,B}_{LM}(f)X^{L}_{\ell\ell^{\prime}}\right]\begin{pmatrix}\ell&\ell^{\prime}&L\\ 2&2&-4\\ \end{pmatrix}\right\},$ (62)
@@
$=2B^{t*}_{\ell}E^{t}_{\ell^{\prime}}I^{t}_{0}(f)\sum_{LM}(-1)^{m+L}\sqrt{\pi}\Braket{\ell(-m)\ell^{\prime}m^{\prime}}{LM}$
×$\qquad\times\left\{\left[-c^{t,I}_{LM}(f)(1-X^{L}_{\ell\ell^{\prime}})+c^{t,V}_{LM}(f)X^{L}_{\ell\ell^{\prime}}\right]\begin{pmatrix}\ell&\ell^{\prime}&L\\ -2&2&0\\ \end{pmatrix}+\left[c^{t,E}_{LM}(f)(1-X^{L}_{\ell\ell^{\prime}})+ic^{t,B}_{LM}(f)X^{L}_{\ell\ell^{\prime}}\right]\begin{pmatrix}\ell&\ell^{\prime}&L\\ 2&2&-4\\ \end{pmatrix}\right\},$ (63)
@@
$=2B^{t*}_{\ell}B^{t}_{\ell^{\prime}}I^{t}_{0}(f)\sum_{LM}(-1)^{m+L}\sqrt{\pi}\Braket{\ell(-m)\ell^{\prime}m^{\prime}}{LM}$
×$\qquad\times\left\{\left[c^{t,I}_{LM}(f)X^{L}_{\ell\ell^{\prime}}-c^{t,V}_{LM}(f)(1-X^{L}_{\ell\ell^{\prime}})\right]\begin{pmatrix}\ell&\ell^{\prime}&L\\ -2&2&0\\ \end{pmatrix}-\left[c^{t,E}_{LM}(f)X^{L}_{\ell\ell^{\prime}}+ic^{t,B}_{LM}(f)(1-X^{L}_{\ell\ell^{\prime}})\right]\begin{pmatrix}\ell&\ell^{\prime}&L\\ 2&2&-4\\ \end{pmatrix}\right\},$ (64)
@@
and the left column in Tables 2 for $F^{L,t,X,EB}_{\ell\ell^{\prime}}$, 3 for $F^{L,t,X,BE}_{\ell\ell^{\prime}}$, and 4 for $F^{L,t,X,BB}_{\ell\ell^{\prime}}$. To get the concrete plots of the ORFs, we similarly take the coordinates given in Eq. (34). Strictly speaking, in that choice of coordinates, $Y^{E}_{(\ell m)\phi}(\hat{n}_{a})$ and $Y^{B}_{(\ell m)\theta}(\hat{n}_{a})$ are not well defined because of $\sin\theta$ in the denominator, which is from the singularity of the coordinates. To avoid this issue, we define $Y^{E}_{(\ell m)\phi}(\hat{n}_{a})$ and $Y^{B}_{(\ell m)\theta}(\hat{n}_{a})$ with $\hat{n}_{a}=\hat{n}_{b}|_{\theta\rightarrow+0}$ in the coordinates. Practically, we take $\hat{n}_{a}=(\sin\theta_{\epsilon},0,\cos\theta_{\epsilon})$ with $\theta_{\epsilon}=10^{-5}$ in the numerical calculation for Figs. 5-8. Then, Figure 5 shows the $\theta$-dependence of $\Gamma^{t,I,EB}_{LM}$. In the figure, we can see $\Gamma^{t,I,EB}_{00}=0$. This physically means that the parity-breaking signal ($EB$-correlation) cannot be produced in the parity-conserving background (isotropic and unpolarized background).